### v1

# Mind-to-Image: Projecting Visual Mental Imagination of the Brain from fMRI

Abstract
The reconstruction of images observed by subjects from fMRI data collected during visual stimuli has made significant strides in the past decade, thanks to the availability of extensive fMRI datasets and advancements in generative models for image generation. However, the application of visual reconstruction has remained limited. Reconstructing visual imagination presents a greater challenge, with potentially revolutionary applications ranging from aiding individuals with disabilities to verifying witness accounts in court. The primary hurdles in this field are the absence of data collection protocols for visual imagery and the lack of datasets on the subject. Traditionally, fMRI-to-image relies on data collected from subjects exposed to visual stimuli, which poses issues for generating visual imagery based on the difference of brain activity between visual stimulation and visual imagery. For the first time, we have compiled a substantial dataset (around 6h of scans) on visual imagery along with a proposed data collection protocol. We then train a modified version of an fMRI-to-image model and demonstrate the feasibility of reconstructing images from two modes of imagination: from memory and from pure imagination. This marks an important step towards creating a technology that allow direct reconstruction of visual imagery.
Machine Learning, ICML
[…]
1 Introduction
Figure 1: Overview of our method. Left: classic fMRI-to-Image approach where a subject sees images in an MRI scanner, bold data are recorded then used to train a model that convert brain data to match the seen images. Right: in our approach we let the subject imagine previously seen images to collect a dataset, then train a model to reconstruct previously seen images from visual imagination brain data. At inference time we collect data where the subject completely imagines new images based on instructions and use the trained model of the first step to generate the imagined images.
In the rapidly evolving field of neuroscience and artificial intelligence, the quest to decode the complex mechanisms of the human brain has led to groundbreaking developments, particularly in the area of functional magnetic resonance imaging (fMRI). Among the myriad applications of fMRI, one of the most intriguing has been the reconstruction of visual experiences. Traditional methods have successfully reconstructed images based on fMRI BOLD (blood-oxygen-level-dependent) data captured while subjects view specific images (Chen et al., 2022, 2023; Scotti et al., 2023). This process involves preprocessing the BOLD data to isolate voxels associated with the brain’s visual areas, generating clip embeddings from these voxels, and then employing image generation algorithms that utilize these embeddings to recreate the observed images. Such methodologies have shown promising results, especially when employing datasets like the NSD (Natural Scenes Dataset) (Allen et al., 2022), a substantial 7T fMRI dataset designed to bridge cognitive neuroscience and artificial intelligence. This dataset, featuring extensive scans from eight subjects, has become a cornerstone for research in this domain.
The intersection of cognitive neuroscience and artificial intelligence holds the potential not only to advance our understanding of brain mechanisms but also to develop practical applications, such as real-time clinical tools and brain-computer interfaces. These applications are predicated on the ability to accurately interpret and reconstruct cognitive states and environmental inputs from brain activity. However, despite the progress in visual reconstruction from observed images, a significant challenge remains in the domain of visual imagination. Visual imagination encompasses both the recall of previously seen images (which we refer to as weak imagination in the paper) and the generation of entirely new images within the mind’s eye (which we refer to as strong imagination). The ability to decode and reconstruct these imagined visuals offers promising possibilities for aiding individuals with disabilities, enhancing creative processes, and even providing novel forms of evidence in legal contexts.
While using models based on reconstruction of images based on visual stimuli to generate images from imagination is theoretically possible, we believe that we need novel approaches to tackle the issue of visualizing imagination in the brain, particularly in data collection and model development. To this end, we introduce a protocol inspired by the NSD dataset but tailored to capture the nuances of visual imagination. Our approach distinguishes between weak and strong imagination, developing specific protocols for each type of visual imagery. We restrict our analysis on two distinct modalities of images: portraits and landscapes, specifically within the theme of surrealism. This choice is not only methodological but also thematic, aligning with the project’s broader artistic aspirations. The intersection of art and science, particularly through the lens of surrealism and its exploration of automatic processes, provides a rich context for our research.
Leveraging the MindEye model (Scotti et al., 2023), a state-of-the-art fMRI-to-image framework, we adapt its architecture to accommodate the broader dimensional complexity inherent in visual imagination data. Our approach is summarised in Fig.1. Our results, while preliminary, indicate the model’s capability to distinguish between and accurately generate portraits and landscapes based on the nature of the imagination. Although the fidelity of the reconstructions to the imagined content varies, our findings underscore the potential of this approach. With further refinement and expansion of the dataset, this research could pave the way for more precise reconstructions of visual imagination, opening new avenues for understanding and exploiting the brain’s creative and cognitive processes.
2 Related work
In the domain of mental imagery, several studies have sought to decode and reconstruct images from the brain’s visual cortex when subjects engage in imagining specific scenes or objects. (Stokes et al., 2009) provided early evidence of the brain’s capacity to activate shape-specific population codes during mental imagery, highlighting the top-down influence of visual imagination on neural activity in the visual cortex. (Reddy et al., 2010) explored the decoding of category information during mental imagery, providing an understanding on how different types of visual information are processed and represented in the brain.
More recently, (Naselaris et al., 2015) developed a voxel-wise encoding model capable of decoding mental images of remembered scenes, demonstrating the relationship between memory, imagination, and visual perception. (VanRullen & Reddy, 2019) employs a combination of variational auto-encoders (VAE) and Generative Adversarial Networks (GAN) to perform robust pairwise decoding of faces with a high degree of accuracy but also to classify gender and decode which face was imagined by subjects. (Goebel et al., 2022) showed the feasibility of reading imagined letter shapes from the mind’s eye using real-time 7 Tesla fMRI, presenting a novel approach to understanding how specific visual information can be decoded from brain activity.
These studies collectively underscore the potential of fMRI data in decoding not just seen images but also those that are imagined or remembered.
3 fMRI data collection protocol for visual imagery
The development of a comprehensive data collection protocol is one the main contributions of our study, aimed at capturing the complex neural correlates of visual imagination. Drawing inspiration from the NSD paper (Allen et al., 2022), our protocol is designed to explore two distinct facets of visual imagination: weak imagination (imagination from memory) and strong imagination (pure imagination). This dual approach enables us to delve into the nuanced ways in which the brain navigates between recalling visual information and generating novel visual constructs.
3.1 Weak imagination: data collection protocol
The protocol for weak imagination aims at recording brain activity corresponding to the recall of previously seen images. In this phase, subjects are first shown an image for 3 seconds, followed by a 1-second rest period. This sequence is repeated with three different images, each selected from our curated set of 600 surrealist portraits and 600 surrealist landscapes (including both real and generated images using Versatile Diffusion (Xu et al., 2023) and Midjourney), amounting to a diverse array of visual stimuli. To engage the brain’s imaginative capacities, the same images are then flashed for 0.1 seconds, prompting the subject to imagine the image for 5 seconds before resting for another second. This process is repeated over a span of 6 hours, encompassing a total of 1200 images. Out of the data collected, 100 data points (comprising brain activity and the corresponding image) are earmarked for validation and evaluation, with the remainder serving as the training set for our fMRI-to-image model. The evaluation focuses on the model’s ability to generate images that not only match the correct modality (portrait or landscape) but also reflect the contents of the original image.
3.2 Strong imagination: data collection protocol
For strong imagination, the process pivots towards the generation of entirely new visual content based on verbal prompts. Subjects are instructed to imagine a portrait or landscape evoking one of ten primal emotions (such as fear, happiness, love, etc.). This task is designed to stimulate the pure imagination process, encouraging the brain to construct novel visual imagery without the reference of previously seen images. Each imagination session lasts for 6 seconds, followed by a 1-second rest, cycling through combinations of landscape/portrait and emotional prompts. This cycle is repeated 10 times. Unlike the data collected for weak imagination, this dataset is solely used during the model’s inference stage. The evaluation here is similar to the weak imagination evaluation: quantitively assess if the model matches the correct modality (portrait or landscape) and qualitatively assess if the contents of the generated image matches the visual imagery of the subject. To ensure that we recall what the subject has imagined, the subject provides an oral description of the visual contents of the imagined image.
3.3 Pre-processing of raw BOLD data and masking of the regions of interest
After the raw BOLD (blood-oxygen-level-dependent) data is collected, we apply a preprocessing routine to provide the model with data that decorrelates all the noise that impair the following steps. We first co-register all BOLD data with a T1 anatomical scan. Then, using a General Linear Model (GLM) with the Glover hemodynamic response function (HRF) (implemented via the Nilearn library (Abraham et al., 2014)), we extracted beta values associated with each voxel for every distinct event, including both the displayed images and the periods of imagination. This allwos us to isolate the specific brain activity corresponding to each event, which is common practice for such tasks.
In the subsequent data pre-processing phase, we identify the brain regions most instrumental to visual imagination and encoding, using a GLM model to manually discern these areas of interest. Unlike the NSD dataset, which predominantly focuses on visual areas, our study necessitated the creation of a custom mask to encapsulate the regions identified as crucial in our specific context of weak and strong imagination. This aforementioned mask was manually curated based on the GLM analysis results, ensuring that only voxels corresponding to our defined regions of interest were included in further analyses. The delineation and analysis of these regions, and how they diverge from the predominantly visual areas highlighted in the NSD study, is studied in the results section.
Finally, we create our training, validation (weak imagination) and test (strong imagination) datasets by flattening the pre-processed betas corresponding the selected voxels and pair them with the images or prompts related to their events.
4 Training the fMRI-to-Image model
We start from the state-of-the-art MindEye model which reconstructs images viewed by subjects from fMRI data by combining two pipelines: a high-level (semantic) pipeline and a low-level (perceptual) pipeline, both grounded in a residual MLP backbone structure. The high-level pipeline maps fMRI voxels to the CLIP Vision Transformer L/14 image space, capturing the semantics of the image. In contrast, the low-level pipeline targets the perceptual fidelity of reconstructed images by mapping voxels to a Variational Autoencoder’s (VAE) image embedding space. They leverage contrastive and mean squared error (MSE) losses for training of its projection and reconstruction capabilities. Techniques such as MixCo are employed to improve classification and retrieval performance by generating synthetic training data.
4.1 Adapting MindEye to larger masks for visual imagery
To accommodate the increased dimensionality inherent in our fMRI BOLD voxel data, which covers extensive brain areas notably larger than only the visual areas of the brain, we adapted MindEye’s architecture. This adjustment aimed to manage the significant growth in input dimensions without exponentially increasing the model’s parameter count. We thus simply refine the architecture of the MLP so that the model efficiently processes the data while maintaining high performance without overburdening computational resources.
4.2 Experiments
Our experimental procedures are as outlined below:
  • We first collect a dataset of weak and strong imagination using the protocol defined in Sec.3, with a total of 6 hours of scans.
  • Then, we train our adapted fMRI-to-Image model using the dataset obtained from the weak imagination data collection protocol. Then, this model is evaluated using the validation set to assess its performance.
  • We finally freeze the parameters of the fMRI-to-Image model. This frozen model is then employed to perform inference on the dataset derived from the strong imagination data collection protocol, with the goal of visualising the purely imagined images.
5 Results
We first analyze the most activate regions of the brain during visual perception, weak imagination and strong imagination.
We then go over the results obtained for training reconstruction fMRI-to-Image models for weak imagination before looking at strong imagination.
5.1 What happens in the brain during visualisation, weak and strong imagination?
In this section we describe the brain regions involved during visual perception, weak imagination, and strong imagination. This is done though the visualisations presented in Fig.2.
Figure 2: Functional MRI results of the three conditions, respectively: simple visual perception (A- in red), weak imagination (B- in pink) and strong imagination (C- in blue) presented through a lower (top row) and right lateral (bottom row) view. Activation maps were generated by comparing the “test” and “rest” conditions, using the same statistical threshold for each test (T-score =5).
The results of the first test (A) show that the brain regions mainly involved in normal visual perception are, as expected, the primary and secondary visual areas in the occipital lobe, the superior parietal lobule and the bottom of the temporal lobe.
Weak imagination (B) involves occipital primary visual areas in a much weaker way. Instead, there is an activation of the basal temporal associative visual areas (fusiform gyrus) involved in mental imagery. The latter are partially overlapped with brain areas involved in visual perception.
Strong imagination (C) involves even more brain areas distinct from visual cortex, with a strong activation in prefrontal regions, involved in the strategy of generating and controlling mental imagery. There is also an incomplete overlap of activations between visual perception and imagination in the temporal lobe, mainly limited to the fusiform gyrus, especially in the right hemisphere, more specialized in the face processing (as this result was obtained during the portrait imagination test).
These results justify the need for a specialised dataset for imagination relative to visual perception, given that the regions most active during the different tests are not quite the same.
5.2 Can we reconstruct images from weak imagination / memory?
We present curated results of the images generated by our model from weak imagination in Fig.3.
Figure 3: Results for weak imagination on the validation dataset.
In the exploration of weak imagination, where subjects were prompted to recall previously seen images, our results highlight the capacity of the model to identify to the original class of the images. Specifically, if a subject recalled a portrait, the reconstructed image is almost always a portrait, and similarly for landscapes. This observation suggests a the ability of the model to grasp the category of the imagined content accurately within the BOLD data. To quantitatively assess this capability, we fine-tune a pretrained ResNet50 model (He et al., 2016) to classify images as either portraits or landscapes using the images in our dataset. We apply this classifier to the reconstructed images from the weak imagination tasks of the validation set. We obtain a classification accuracy of 91% in correctly identifying the reconstructed images as belonging to their original categories of either portraits or landscapes. This level of accuracy demonstrates the model’s ability in capturing the category information within the neural activity during weak imagination.
While the model demonstrated a robust ability to regenerate the broad category of the original images, the fidelity of content reconstruction within these images presented a more varied outcome. Some reconstructions were close to the remembered images. However, others diverged more significantly, indicating a variability in the model’s ability to precisely capture and reconstruct the nuanced details of each imagined scene. This variance in reconstruction fidelity can be attributed, in part, to the limitations imposed by the dataset’s size and the duration of the fMRI scanning sessions. With only 6 hours of scans and 1200 images constituting the dataset, the goal for training a model from scratch to achieve perfect reconstruction is out of reach. Nevertheless, these initial results are promising, suggesting a capacity for accurate content reproduction. It is reasonable to anticipate that with an expanded dataset, the fidelity of the reconstructed images to their original counterparts would see significant improvement, enhancing both the precision and consistency of the content reproduction in future iterations.
5.3 Can we reconstruct images from strong imagination / pure imagination?
We present curated results of the images generated by our model from strong imagination in Fig.4.
Figure 4: Results from strong imagination.
In the phase dedicated to strong imagination, we employed the previously trained and then frozen fMRI-to-Image model to process BOLD data obtained while subjects engaged in pure imagination tasks, such as imagining a portrait or landscape embodying a concept like optimism. This procedure is effectively transfer learning, leveraging a model trained exclusively on weak imagination data to interpret and generate visual outputs from the distinct context of strong imagination, without fine-tuning. In any case, we do not have tangible images to serve as direct training data for strong imagination, since these images exist solely within the mind of the subject. Hence we employ transfer learning from the weak imagination process to generate from strong imagination. Remarkably, the application of the fine-tuned ResNet50 classifier to the generated images resulted in an 88% accuracy rate in categorizing the images according to the instructions provided to the subjects. This accuracy in category generation demonstrates the effectiveness of transfer learning from weak to strong imagination, which represent our main contribution.
Determining the content accuracy of images generated from strong imagination poses a unique challenge, as the ’ground truth’ images solely exist in the subject’s mental imagery, inaccessible for direct comparison. To navigate this, we based our evaluation on subjects’ descriptions of their imagined scenes, collected during or following the scanning sessions. The variability in the reconstructed content ranges from highly accurate renditions of specific imagined elements, like a tree, dark-haired woman, or flowers, to more abstract interpretations or plain errors. This also reflects the subjective nature of this task. Given the reliance on subjective descriptions and the inherent open-endedness of strong imagination, the results should be interpreted with caution. While it’s challenging to draw definitive conclusions about the model’s capacity for precise content reconstruction in this context, the instances of recognisable elements emerging in the reconstructions hint at the potential depths of accuracy achievable with increased data volume and further model refinement.
6 Discussion
6.1 Ethical considerations
As we explore further into the capabilities of decoding and reconstructing visual content from brain activity, we get closer to the boundaries of personal thought and imagination, a concept known as ”mind privacy”. The ability to visualise an individual’s mental imagery raises significant questions about consent, mental autonomy, and the potential for misuse of such technologies. On the one hand, such technology provides potential for good, opening avenues for novel communication methods for those unable to speak or write, enhancing creative expression, and deepening our understanding of the human brain and its processes. On the other hand, it introduces the possibility of invasive breaches of personal thought, where the most private mental images and ideas could be exposed without consent. Navigating these waters requires a careful, deliberate approach to ensure that the development and application of these technologies are governed by ethical principles that prioritise individual privacy and autonomy. Currently, the barriers to entry, including the need for explicit consent and the use of expensive and sophisticated MRI machinery, provide a layer of protection.
6.2 Future work
Future directions for this line of research include the acquisition of more extensive data to improve the accuracy and specificity of content match in the reconstructed images. A refined experimental protocol for both weak and strong imagination tasks could lead to more nuanced insights into the brain’s imaging processes, while improved evaluation methods for strong imagination would enhance our understanding of the model’s effectiveness in capturing the depth and detail of imagined content. Additionally, the question of whether MRI technology can be replaced or complemented by EEG is promising. Recent studies suggest the feasibility of reconstructing visual stimuli from brain activity using EEG (Benchetrit et al., 2023), which offers a more accessible, and cost-effective alternative to MRI.
7 Conclusion
In our work, we introduced and applied two novel data collection protocols: weak and strong imagination. These protocols were designed to study visual imagery and create training datasets for fMRI-to-Images models. We implemented them during a 6-hour scanning session to create a comprehensive dataset. This dataset then served as the basis for training our fMRI-to-image model. Initially, the model was trained using data from the weak imagination protocol, which involved reconstructing images that the subject had previously seen. Subsequently, the same model, now frozen, was applied to data from the strong imagination protocol, where the subject was asked to imagine new images based on a written instruction. Our study demonstrates that the model successfully distinguishes between broad categories such as portraits and landscapes, effectively grasping the category of the content imagined by the subject. Accurately capturing the detailed contents of these imagined images was partly successful, but proved to be more challenging. This research provides a promising path towards the generation of visual representations directly from human thought.
Acknowledgements
The author acknowledges the support of the French Agence Nationale de la Recherche (ANR) under reference ANR-23-CE23-0023.